Use the cross product to find the sine of the angle between the vectors u = (2,8,-2) and v = (2,8,2) .
1
Expert's answer
2011-11-24T09:12:04-0500
u = (2,8,-2) and v = (2,8,2)
Let's use the formula:
|uXv| =
|u|*|v|*sinA.
uXv = (u2v3 - u3v2)*i + (u3v1 - u1v3)*j + (u1v2 - u2v1)*k
=
(8*2 + 2*8)*i + (-2*2 - 2*2)*j + (2*8 - 8*2)*k = 32i - 8j + 0k,
|uXv| =
sqrt(32^2 + 8^2) = sqrt(1088),
|u|*|v| = sqrt(2^2 + 8^2 + 2^2)*sqrt(2^2 + 8^2
+ 2^2) = 72.

So, sinA = sqrt(1088) / 72 = 0.458.